Fermi National Laboratory


Accelerator Update

Friday June 25
The day shift began with Operations monitoring store 3592, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MTest, and MCenter experiments taking beam.

At 9:57 AM, a Linac RF station (LRF5) tripped off due to a breaker trip. Operators had it back on within five minutes.

At 11:23 AM, operators turned off the Booster RF system to investigate a water flow problem. At 11:57 AM, operators bypassed a RF station (BRF15) for work. Pbar resumed stacking.

Operations reported at 12:09 PM that the Tevatron (TeV) seismic sensors noted some activity just as monitors noted some minor beam problems.

Operations took the Booster offline for a time while technicians reconnected BRF15 back into the system. Pbar resumed stacking and MiniBooNE resumed taking beam at 12:19 PM.

At 1:26 PM, a power glitch hit the Fixed Target areas and CHL. The high voltage expert reported that feeder 31 glitched. Operations warned the Collider Detector at Fermilab (CDF) and D-Zero (D0) that with CHL problems, the TeV cryo systems might trip and cause the store to abort. The experiments turned off their detectors.

A Main Injector (MI) expert began a slipstacking study at 2:14 PM. During this study, Pbar saw a RECORD amount of beam (6.86E12) on their antiproton target.

At 3:09 PM, Operations reported that the CHL cryo system problems had recovered.

The slipstacking study ended at Pbar resumed normal stacking at 5:02 PM.

At 9:32 PM, the TeV store 3592 aborted due to a quench at sector B4. The quench was due to a cryo system crate power supply failing. Operators replaced the supply.

Operators prepared the CDF and D0 collision halls for an access at 10:47 PM.

At 11:05 PM, the MiniBooNE horn tripped off. It reset okay but then tripped off again five minutes later. Operators reduced the horn's rate so that it would stay on and called an expert.

Saturday June 26
The midnight shift began with the TeV recovering from a quench, with CDF and D0 in access, with Pbar stacking, and with MiniBooNE taking beam.

Operations put the TeV through a dry squeeze at 2:24 AM and then into shot setup at 2:56 AM.

Operations established store 3594 at 5:30 AM with an initial luminosity of 81.83E30. Pbar resumed stacking.

Operations stopped stacking and setup the Recycler for stashing antiprotons at 7:48 AM.

The Recycler stash ended 9:23 AM.

Pbar experts began a Zero-stack study at 9:55 AM. This study halted all Switchyard (SY) beam.

A Linac RF station (KRF5) tripped off at 9:59 AM. It tripped off five more times while operators tried to turn it back on. Operators managed to get the station on and back up six minutes later.

Operators began preparing MCenter for beam 11:16 AM.

Pbar experts completed their Zero-stack study at 1:06 PM. Pbar resumed stacking.

Operations established beam to MCenter at 1:30 PM.

Sunday June 27
The midnight shift began with Operations monitoring store 3594 and the Recycler stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 1:22 AM, KRF5 tripped off. It reset okay.

At 1:56 AM, a Booster RF station (BRF4) tripped off. It reset okay.

MCenter accessed their collision hall at 3:10 AM and then dropped their interlock system. Operators search and secured their enclosure and had them taking beam by 4:32 AM.

Operations terminated store 3594 at 9:13 AM.

Operations made CDF and D0 ready for a forty-five minute access at 9:30 AM.

At 11:31 AM, two Booster RF stations (BRF7 & 8) tripped off due to a vacuum high-pressure indication. Operators discovered a vacuum pump off.

Operations put the TeV into shot setup at 11:32 AM and established store 3596 at 2:41 PM. The store's initial luminosity was 73.85E30. Pbar resumed stacking and MCenter resumed taking beam.

Operations setup the Recycler to stash antiprotons at 5 PM.

Pbar resumed stacking and MCenter resumed taking beam at 6 PM.

At 9:05 PM, a Booster RF station (BRF1) tripped off for the first of seven times in a little over three hours.

Operators began sending the MTest experiment beam around 9:15 PM.

Monday June 28
The midnight shift began with Operations monitoring store 3596 and the stash, with Pbar stacking, and with MiniBooNE, MCenter, and MTest taking beam.

MTest finished using beam 12:20 AM.

At 1:12 AM, the TeV suffered a sector E1 quench due to an earthquake centered eight miles north west of Ottawa, Illinois.

Operations prepared D0 for an access at 1:48 AM.

Operations put the TeV into shot setup at 5:09 AM.

Around 5 AM, Operations noted more TeV seismic activity. A 6.5 level earthquake hit Alaska in the Queen Charlotte Islands region.

Here is what the quake looked like to operators.

This quake would probably have caused a quench if there was a store in the TeV.

Operations established store 3597 at 7:53 AM with an initial luminosity of 63.71E30.

Plans for Today
The plans for today include Recycler stashing antiprotons later this morning followed by Pbar experts conducting a Zero-stack study. Tomorrow, Operations and experts will establish a Mixed Pbar store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/28/2004   email Fermilab